# NeuroTree: Hierarchical Functional Brain Pathway Decoding for Mental Health Disorders

### v2

@@ Abstract
Analyzing functional brain networks using functional magnetic resonance imaging (fMRI) is crucial for understanding psychiatric disorders and addictive behaviors. While existing fMRI-based graph convolutional networks (GCNs) show considerable promise for feature extraction, they often fall short in characterizing complex relationships between brain regions and demographic factors and accounting for interpretable variables linked to psychiatric conditions. We propose NeuroTree to overcome these limitations, integrating a $k$-hop AGE-GCN with neural ordinary differential equations (ODEs). This framework leverages an attention mechanism to optimize functional connectivity (FC), thereby enhancing dynamic FC feature learning for brain disease classification. Furthermore, NeuroTree effectively decodes fMRI network features into tree structures, which improves the capture of high-order brain regional pathway features and enables the identification of hierarchical neural behavioral patterns essential for understanding disease-related brain subnetworks. Our empirical evaluations demonstrate that NeuroTree achieves state-of-the-art performance across two distinct mental disorder datasets and provides valuable insights into age-related deterioration patterns. These findings underscore the model’s efficacy in predicting psychiatric disorders and elucidating their underlying neural mechanisms.
{IEEEkeywords}
fMRI, graph neural networks, mental health disorders, neural ODE, brain tree
@@ 1 Introduction
In recent years, neuropsychiatric disorders and addiction have emerged as substantial public health challenges, with their impact on brain function becoming a critical research focus in neuroscience [1, 2, 3]. Substance use disorder (SUD) is widely studied in the field of mental disorders, enabling the investigation of structural brain alterations among SUD users across various factors (e.g., age or gender) [4, 5]. The development of functional magnetic resonance imaging (fMRI), which measures blood oxygen level-dependent (BOLD) signals, has provided unprecedented opportunities to analyze patterns of brain network activity in mental disorders [6, 7, 4].
@@
Moreover, research on the clustering of connectivity in brain regions using fMRI has demonstrated that the brain does not operate statically but instead exhibits a dynamic, multi-level hierarchical organization [8, 9]. In particular, hierarchical network analysis has been shown to effectively distinguish these abnormal patterns and identify brain regions associated with specific pathological mechanisms in psychiatric disorders [10, 11].
@@
Currently, traditional brain network analysis has successfully represented regions of interest (ROIs) as nodes with functional connectivity (FC) as edges [12, 13, 14]. Recent research on graph neural networks (GNNs) has shown promising results, demonstrating strong predictive capabilities for neuropsychiatric disorders by learning patterns across brain networks [15, 16].
@@
Differential equations can capture the dynamic processes underlying changes in brain activity over time. Several studies have employed conventional network-based and dynamic causal modeling (DCM) approaches to capture neural-hemodynamic relationships [17, 18]. CortexODE further advances these approaches by combining ordinary differential equations (ODEs) with deep learning methods to model MRI images as point trajectories, thereby enabling the reconstruction of complex cortical surfaces [19]. Moreover, STE-ODE [20] combines the advantages of ODEs with continuous-time fMRI to model dynamic brain networks while leveraging GNN architectures for embedding learning. However, these approaches have several limitations. First, although graph convolutional networks (GCNs) demonstrate excellent node feature learning capabilities in graph structures, they are generally constrained by limited interpretability—particularly regarding the explicability of auxiliary factors in mental disorders. Second, graph-based networks for human brain functional connectivity are restricted to feature learning from dynamic graph sequences, overlooking the brain’s hierarchical clustering of network activity. Consequently, there is a lack of effective methodologies for predicting functional brain alterations across different age stages in psychiatric disorders.
@@
To address these limitations, we leverage the properties of ODEs for fMRI modeling like STE-ODE’s continuous-time dynamics methods. In this study, we propose NeuroTree, a framework designed to characterize brain network patterns through hierarchical, tree-structured pathways. Drawing from neuroscience research indicating that psychiatric disorders affect brain regions in hierarchical patterns, we represent the fMRI network as tree structures. This approach decodes brain network features and facilitates visualization of disorder propagation through brain pathways. Additionally, we develop a $k$-hop ODE-GCN to capture complex interactions between adjacent and distant brain regions and to measure the convergence of various mental disorder subtypes. We incorporate demographic features like age into node representations and message passing to enhance interpretability. Moreover, we implement contrastive masking optimization to identify key functional connectivity patterns. Our main contributions are as follows:
  • We integrate interpretable variables into static and dynamic causal graph convolutional modeling through $k$-hop convolution.
@@
  • We investigate two types of brain disorder datasets using fMRI brain networks, constructing disease-specific brain trees while employing trainable $k$-hop graph convolutions to identify optimal high-order tree paths.
@@
  • Our research not only achieves state-of-the-art graph classification performance but also effectively predicts and interprets FC brain age change patterns associated with addiction and mental disorders.
Figure 1: Overview of NeuroTree framework.
@@ 2 Related Work
2.1 High-Order GCNs for Brain Disease Classification.
In the auxiliary diagnosis of brain diseases, feature extraction and classification methods based on functional connectivity networks (FCNs) have received widespread attention. Early research primarily focused on low-order functional connectivity (LOFC), constructing networks by calculating temporal correlations between brain regions [21, 22]. However, considering only pairwise connections between nodes may overlook information about higher-level interactions among brain regions. Therefore, researchers introduced high-order functional connectivity networks (HOFC), which capture more complex brain network characteristics by calculating correlations between functional connectivity distribution sequences [23, 24]. Moreover, recent work has shown that traditional first-order GCNs rely on direct neighbor information but struggle to model the long-range dependencies critical for understanding brain functional networks. For instance, one study proposed a multi-channel fusion GCN (MFGCN) that integrates both topographical and high-order functional connectivity for depressive disorder (MDD) classification [25].
The design of higher-order neighbor matrices—exemplified by HWGCN—defines $k$-hop neighbor matrices based on shortest path distances to learn weighted matrices, effectively integrating first-order and higher-order neighborhood information [26]. In fMRI studies of brain disorders, learned masks have been utilized to construct hyperedges and compute weights for each hyperedge in hypergraph construction, revealing maximum information minimum redundancy (MIMR) higher-order relationships among brain regions [27].
2.2 Hierarchical and Path-Aware Learning in GNNs.
Some studies have found that noise connectivity issues in brain networks and the difficulty of stacking deep layers in traditional GCN models lead to over-smoothing problems. Researchers have proposed hierarchical GCNs to capture richer feature representations [28]. For example, HFBN-GCN designed a coarse-to-fine hierarchical brain feature extraction process with two branches: one focusing on local information transmission within the brain’s hierarchical structure and another processing the entire brain network to preserve the original information flow [29].
Unlike hierarchical GCN approaches, GCKN computes node distances using kernel functions along vertex-connected paths [30]. Additionally, PathNNs propose an end-to-end approach for extracting multiple types of paths and updating node representations by encoding paths originating from nodes into fixed-dimensional vectors through LSTM-based path aggregation [31]. The HEmoN model [32] has provided preliminary insights into human emotions through tree path analysis; however, it lacks interpretable representations of higher-order brain pathways. In this study, we propose a weighted higher-order brain tree in a hierarchical pathway framework that enhances pathway representation and effectively explains the predictions of age-related changes in FC networks associated with mental disorders across different age groups.
@@ 3 Preliminaries
3.1 Problem Formulation and Notations.
In this section, we formulate the mental disorder classification problem as a graph classification task, using fMRI networks constructed for each patient. We introduce the temporal brain activity measurements by giving $N$ patients with fMRI data. Let $X_{i}(t)\in\mathbb{R}^{v}$ denote the blood oxygen level-dependent (BOLD) signals at time point $t$ for subject $i$, where $v$ represents the number of regions of interest (ROIs). The complete time series for each subject can be represented as $X_{i}\in\mathbb{R}^{v\times T}$, where $T$ is the length of the time series. The foundation of brain network analysis lies in representing these temporal patterns through graph structures. For each subject, we construct a functional connectivity (FC) graph $G(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of vertices (i.e., brain regions) and $\mathcal{E}$ denotes the set of edges (i.e., functional connections). The functional connectivity between regions is typically computed using Pearson or partial correlation methods. Based on these temporal correlation patterns, we derive both a static adjacency matrix $A^{s}$ from the complete time sequence and time-varying dynamic FC matrices $A^{d}(t)$. The static matrices capture averaged temporal dependencies, while the dynamic matrices preserve the temporal evolution of neural interactions [6, 7].
3.2 Graph Convolutional Networks.
@@
where $A$ is the adjacency matrix, $D$ is the degree matrix, $H^{(l)}$ represents the node features at layer $l$, and $W^{(l)}$ contains learnable parameters. However, this formulation is limited to first-order neighborhoods and may not effectively capture the complex interactions among brain region [33].
3.3 Neural ODEs for Brain Dynamics.
Recent work has demonstrated the effectiveness of ordinary differential equations (ODEs) in modeling continuous-time neural dynamics [20]. The evolution of brain states can be described by the equation $\frac{dX(t)}{dt}=f(X(t),t)$, where $X(t)$ represents the brain state at time $t$ and $f(\cdot)$ is a neural network that learns the rules governing state transitions. This framework provides a natural way to model the continuous nature of brain dynamics while maintaining biological plausibility [17].
4 The Proposed NeuroTree
The proposed NeuroTree framework, illustrated in Fig. 1, constructs dynamic FC graphs from subjects’ time-series data. It encodes dynamic graph fMRI features using a $k$-hop ODE-GCN for graph embedding and FC strength computation. Subsequently, a CMFS optimization is performed to adjust the FC strength between similar and dissimilar brain regions, followed by tree pruning to create anatomically meaningful hierarchical brain network structures. The resulting brain tree structures are then employed to interpret differences between healthy controls and disease cohorts. Finally, NeuroTree predicts functional age-related changes and compares these with chronological age.
@@ 4.1 $K$-hop ODE-Based Graph Embedding
The ODE effectively describes the dynamic changes in effective connectivity between different ROI[34]. Inspired by existing studies [35, 36, 20] that utilize ODEs to model blood-oxygen-level-dependent signals or biomarkers in Alzheimer’s disease (AD) research, we extend the ODE modeling framework by incorporating a scalar parameter, $\theta\in\mathbb{R}^{+}$, representing the chronological age of subjects. This enhancement aims to improve the prediction and interpretability of mental disorders in fMRI analysis.
@@
where the matrix $C$ encodes the reception of external stimuli $u(t)$ and parameters $\eta$ and $\rho$. To simplify the analysis, we assume that $Cu(t)=0$, indicating the absence of additional external stimuli. By discretizing Eq. 3, we derive the effective connectivity matrix $A^{d}(t)$. Its discrete representation is given by:
@@
where $\varphi=\frac{1}{\eta}$ and $\rho$ are a scaling factors ranging from 0 to 1. To address the limitations of first-order feature convolution filters in traditional GCNs, we propose an extended spatial feature filter that operates over a $k$-hop connectivity operator. This approach enables the spatial $k$-order brain network representation to be learned via a generalized graph convolution.
@@
where $\Gamma$ is a learnable parameter, the $k$-hop connectivity operator retains directional information, as evidenced by the fact that $\hat{A}_{k}(t)\neq(\hat{A}_{k}(t))^{T}$ when $\lambda\neq\frac{1}{2}$. This result is derived using the binomial theorem, as detailed in Appendix 12.
Theorem I. Suppose the $l^{2}$-norm of the $k$-hop connectivity adjacency operator $\|\hat{A}_{k}(t)\|_{2}$ can be derived from Eq. (7). Then, as the $k$-hop approaches infinity ($k\to\infty$ ), the $\|\hat{A}_{k}(t)\|_{2}$ is bounded by the following inequality:
@@
The detailed proof of Eq. 8 is provided in Appendix 13. To ensure numerical stability when computing powers of $\hat{A}_{k}(t)$, we establish the bounded condition: $\|\Phi_{k}(t)\|_{2}\leq 1$, which holds for all values of $k$ and $t$. In proposition I, we demonstrate that $\Phi_{k}(t)$ exhibits smooth temporal evolution of the graph structure while maintaining a well-defined upper bound.
Proposition I. Let $\Phi_{k}(t)$ be the normalized adjacency operator that satisfies the Lipschitz condition [37] $\|\Phi_{k}(t_{1})-\Phi_{k}(t_{2})\|_{2}\leq L\|t_{1}-t_{2}\|_{2}$ for some constant $L\geq 0$. Then, for a Lipschitz continuous activation function $\sigma$ and bounded weights $W_{k}^{(l)}$, the $k$-hop ODE-GCN admits a unique solution. The proof from Picard’s existence theorem can be found in Appendix 14.
Theorem II. [20] Consider a series of brain networks $\{\mathcal{G}(t;\theta)\}_{t=0}^{T}$ parameterized by an age variable $\theta$. Let $Z(t)$ denote the embedding (or feature representation) of the STE-ODE network at time $t$. The continuous temporal evolution of the brain network can be expressed through the following age-modulated ODE-based graph convolution representation:
@@
with $H^{(l)}(t,\theta)$ representing the age-modulated layer-wise transformation at time $t$. Using a first-order approximation, the temporal dynamics described in Eq. 6 and theorem theorem II can be discretized into the following form:
@@
$=Z(t)+\Delta t\sum_{l=1}^{L}\sigma\Big{(}\sum_{k=0}^{K-1}\Phi_{k}(t)X_{\theta}^{(l)}(t)W_{k}^{(l)}\Big{)}.$
Here, $X_{\theta}^{(l)}(t)=\Lambda\theta H^{(l)}(t)$ denotes the age-modulated features at layer $l$, where $\Lambda$ is a trainable parameter. In this context, $\Phi_{k}(t)$ represents the polynomial filter coefficients at time $t$, $H^{(l)}(t)$ is the feature matrix at layer $l$, and $W_{k}^{(l)}$ are the learnable weights associated with the $k$-th hop filter at layer $l$.
For $t\geq 1$, the temporal graph embedding update is expressed as:
@@
4.2 Attention-Based FC Strength Enhancing
To enhance the FC between neighborhood nodes representations, we further implement an attention mechanism that computes attention for each timestamp $t\in\{1,2,...,T\}$ in the temporal networks $A^{d}(t)$ for each node $i$. Specifically, we encode the $t$th adjacency matrix $A^{d}$ into a features matrix: $\mathcal{H}=A^{d}W$ and then map it into three embeddings: the query $\mathcal{H}W^{Q}_{i}$, key $\mathcal{H}W^{K}_{i}$, and value $\mathcal{H}W^{V}$ matrices, where $W^{Q}$,$W^{K}$, and $W^{V}$ $\in\mathbb{R}^{d\times d_{k}}$ are learnable matrices. A single attention head is computed as:
$\text{head}_{i}=\text{softmax}\left(\frac{\mathcal{H}W_{i}^{Q}(\mathcal{H}W_{i}^{K})^{T}}{\sqrt{d}_{k}}\right)\mathcal{H}W_{i}^{V}$ (13)
where $d_{k}$ represents the dimensionality of each attention head. Next, we concatenate the outputs of all $h$ heads to form the multi-head attention (MHA):
$\mathcal{H}^{\text{MHA}}=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W^{O}$ (14)
where $W^{O}\in\mathbb{R}^{hd_{k}\times d}$ is the projection matrix. For simplicity, we use feature-wise averaging over the $d$-dimensional FC output:
$C_{i}(t)=\frac{1}{d}\sum_{k=1}^{d}\mathcal{H}_{i,k}^{\text{MHA}}$ (15)
where $\mathcal{H}^{\text{MHA}}_{i,k}$ denotes the $k$-th dimension of the $i$-th node’s representation in $\mathcal{H}^{\text{MHA}}$.

### v3

@@ Abstract
Mental disorders are among the most widespread diseases globally. Analyzing functional brain networks through functional magnetic resonance imaging (fMRI) is crucial for understanding mental disorder behaviors. Although existing fMRI-based graph neural networks (GNNs) have demonstrated significant potential in brain network feature extraction, they often fail to characterize complex relationships between brain regions and demographic information in mental disorders. To overcome these limitations, we propose a learnable NeuroTree framework that integrates a $k$-hop AGE-GCN with neural ordinary differential equations (ODEs) and contrastive masked functional connectivity (CMFC) to enhance similarities and dissimilarities of brain region distance. Furthermore, NeuroTree effectively decodes fMRI network features into tree structures, which improves the capture of high-order brain regional pathway features and enables the identification of hierarchical neural behavioral patterns essential for understanding disease-related brain subnetworks. Our empirical evaluations demonstrate that NeuroTree achieves state-of-the-art performance across two distinct mental disorder datasets. It provides valuable insights into age-related deterioration patterns, elucidating their underlying neural mechanisms. The code and datasets are available at https://github.com/Ding1119/NeuroTree.
Machine Learning, ICML
Figure 1: Overview of NeuroTree framework.
@@ 1 Introduction
In recent years, neuropsychiatric disorders and addiction have emerged as substantial public health challenges, with their impact on brain function becoming a critical research focus in neuroscience (Sahakian et al., 2015; Hollander et al., 2020; Borumandnia et al., 2022). Substance use disorder (SUD) is widely studied in the field of mental disorders, enabling the investigation of structural brain alterations among SUD users across various factors (e.g., age or gender) (Niklason et al., 2022; Ding et al., 2024b). The development of functional magnetic resonance imaging (fMRI), which measures blood oxygen level-dependent (BOLD) signals, has provided a technique to analyze patterns of brain network activity in mental disorders (Zhang et al., 2016; Li et al., 2022a; Niklason et al., 2022).
@@
Moreover, research on the clustering of connectivity in brain regions using fMRI has demonstrated that the brain does not operate statically but instead exhibits a dynamic, multi-level hierarchical organization (Meunier et al., 2009; Betzel et al., 2023). In particular, hierarchical network analysis has been shown to effectively distinguish these abnormal patterns and identify brain regions associated with specific pathological mechanisms in psychiatric disorders (Boisvert et al., 2024; Segal et al., 2023).
@@
Currently, traditional brain network analysis has successfully represented regions of interest (ROIs) as nodes with functional connectivity (FC) as edges (Kulkarni et al., 2023; Xu et al., 2024, 2016). Specifically, research on graph neural networks (GNNs) has shown promising results, demonstrating strong predictive capabilities for neuropsychiatric disorders by learning patterns across brain networks (Tong et al., 2023; Kazi et al., 2019).
@@
Differential equations can capture the dynamic processes underlying changes in brain activity over time. CortexODE further advances these approaches by combining ordinary differential equations (ODEs) with deep learning methods to model MRI images as point trajectories, thereby enabling the reconstruction of complex cortical surfaces (Ma et al., 2022). Moreover, STE-ODE (Tang et al., 2024) combines the advantages of ODEs with continuous-time fMRI and structural MRI to model dynamic brain networks while leveraging GNN architectures for embedding learning. However, these approaches have several limitations. First, although GNNs demonstrate excellent node feature learning capabilities in graph structures, they are generally constrained by limited interpretability—particularly regarding the explicability of auxiliary factors in mental disorders. Second, graph-based networks for human brain FC are restricted to feature learning from dynamic graph sequences, overlooking the brain’s hierarchical clustering of network activity. Consequently, there is a lack of effective methodologies for predicting functional brain alterations across different age stages in psychiatric disorders.
@@
To address these limitations, we leverage the properties of ODEs for fMRI modeling similar to STE-ODE and extend continuous-time dynamics methods. In this study, we propose NeuroTree, a framework designed to characterize brain network patterns through hierarchical analysis and represent the fMRI network as tree-structured pathways. This approach decodes brain network features and facilitates visualization of disorder propagation through brain pathways. Additionally, we develop a $k$-hop ODE-based AGE-GCN to capture complex interactions between adjacent and distant brain regions and to measure the convergence of various mental disorder subtypes. We incorporate demographic features like age into model learning and message passing to enhance interpretability. Moreover, we implement contrastive masking optimization to identify key FC patterns. Our main contributions are as follows:
  • We integrate demographic information into Neural ODEs for modeling both static and dynamic brain networks via $k$-hop graph convolution.
@@
  • We investigate two types of brain disorder datasets to decode fMRI signals, constructing disease-specific brain trees while employing trainable framework to identify hierarchical functional subnetwork regions.
@@
  • Our research not only achieves state-of-the-art graph classification performance but also effectively interprets how addiction and schizophrenia disorders cause changes in FC related to brain age.
@@ 2 Related Work
Neural ODEs for Dynamic BOLD fMRI Signal Modeling.
The BOLD fMRI signal is an indirect and delayed reflection of neural activity. Traditional approaches like DCM use ordinary differential equations (ODEs) to infer connectivity, but they are computationally intensive and rely on prior assumptions  (Cao et al., 2019). Recent work has demonstrated the effectiveness of ODEs in modeling continuous-time neural dynamics (Han et al., 2024; Tang et al., 2024). The evolution of brain states can be described by the equation $\frac{dX(t)}{dt}=f(X(t),t)$, where $X(t)$ represents the brain state at time $t$, and $f(\cdot)$ is a neural network that learns the rules governing state transitions. This framework provides a natural way to model the continuous nature of brain dynamics while maintaining biological plausibility (Havlicek et al., 2015).
Path-Aware Learning in GNNs.
GCKN computes node distances using kernel functions along vertex-connected paths (Chen et al., 2020). Additionally, PathNNs enhance the expressive power of GNNs by aggregating information from multiple paths between nodes to improve performance across multiple graph classification and regression tasks and capturing complex graph structures (Michel et al., 2023). The HEmoN model (Huang et al., 2025) has provided preliminary insights into human emotions through tree path analysis; however, it lacks aggregation of higher-order brain pathways in complex brain region interactions, as well as consideration of different demographic factors between individuals. In this study, we propose a weighted higher-order brain tree pathway framework that enhances pathway representation and effectively explains the predictions of changes in FC networks associated with mental disorders across different age groups.
3 The Proposed Architecture
The proposed NeuroTree framework, illustrated in Fig. 1, constructs dynamic FC graphs from subjects’ time-series data. It encodes dynamic graph fMRI features using a $k$-hop ODE-Based AGE-GCN incorporate demographics for graph embedding learning for downstream tasks (e.g., graph classification, brain age prediction). Subsequently, a CMFC optimization is performed to adjust the FC strength between similar and dissimilar brain regions distance. Then, we build the hierarchical brain tree by reranking predicted brain regions of interest (ROIs) scores and high-order tree path weights, followed by tree pruning to create anatomically meaningful hierarchical brain network structures. The resulting brain tree structures are then employed to interpret disease cohorts, where functional networks primarily affecting mental disorders manifest at different tree levels.
@@ 4 Preliminaries
Problem Formulation and Notations. In this section, we formulate the classification of mental disorders using both static and sequential segmented instance brain networks constructed from fMRI as a supervised graph classification task for each patient. Let temporal brain activity measurements from $N$ patients, $X_{i}\in\mathbb{R}^{v\times T}$, represent the complete BOLD fMRI signals with $v$ regions of interest (ROIs) and a total length of $T$ time series. We denote $X_{i}(t)\in\mathbb{R}^{v\times T^{\prime}}$ is segmented time points with dimension $T^{\prime}$ for the $i$th subject. For each subject, we construct a static or dynamic FC graph $G(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of vertices (i.e., brain regions) and $\mathcal{E}$ denotes the set of edges (i.e., functional connections). The FC between regions is typically computed using Pearson correlation methods. Based on these temporal correlation patterns, we derive both a static adjacency matrix $A^{s}\in\mathbb{R}^{v\times v}$ from the complete time sequence and time-varying dynamic FC matrices $A^{d}(t)\in\mathbb{R}^{v\times v}$. The static matrices capture averaged temporal dependencies, while the dynamic matrices preserve the temporal evolution of neural interactions (Zhang et al., 2016; Li et al., 2022a).
4.1 Graph Convolutional Networks
@@
where $A$ is the adjacency matrix, $D$ is the degree matrix, $H^{(l)}$ represents the node features at layer $l$, and $W^{(l)}$ contains learnable parameters. However, this formulation is limited to first-order neighborhoods and may not effectively capture the complex interactions among brain region (Veličković et al., 2017).
@@ 4.2 $K$-hop ODE-Based Graph Embedding
The ODE effectively captures the long-term dependencies of dynamic changes in effective connectivity between different ROIs (Sanchez-Romero et al., 2019). Inspired by existing studies (Cao et al., 2019; Wen et al., 2024; Tang et al., 2024) that utilize ODEs to model BOLD signals or biomarkers in Alzheimer’s disease (AD) research, we extend the ODE modeling framework by incorporating a scalar parameter, $\theta\in\mathbb{R}^{+}$ (i.e., age) , representing the chronological age of subjects. This enhancement aims to improve the prediction and interpretability of mental disorders in fMRI analysis.
@@
where the matrix $C\in\mathbb{R}^{v\times v}$ encodes the reception of external stimuli $u(t)\in\mathbb{R}^{v}$ and parameters $\eta$, $\rho\in\mathbb{R}$. To simplify the analysis, we assume that $Cu(t)=0$, indicating the absence of additional external stimuli due to resting-state fMRI task. By discretizing Eq. 3, we can derive the effective connectivity matrix $A^{d}(t)$. Its discrete representation is given by:
@@
where $\varphi=\frac{1}{\eta}$ and $\rho$ are scale factors ranging from 0 to 1. To address the limitations of first-order feature convolution filters in traditional GCNs, we propose an extended spatial feature filter that operates over a $k$-hop connectivity operator. This approach enables the spatial $k$-order brain network representation to be learned via a generalized graph convolution, as defined in definition 4.1.
Definition 4.1 ($k$-hop graph convolution).
The spatial $k$-order brain network representation can be learned through a generalized graph convolution as follows:
@@
where $\Gamma$ is a learnable parameter, the $k$-hop connectivity operator retains directional information, as evidenced by the fact that $\hat{A}_{k}(t)\neq(\hat{A}_{k}(t))^{T}$ when $\lambda\neq\frac{1}{2}$. This result is derived using the binomial theorem, as detailed in Appendix A. Intuitively, Eq. 7 considers that $A^{s}$ provides connectivity information for the entire static fMRI sequence, while $A^{d}$ can regulate the bidirectional combination of dynamic fMRI in the brain through the parameter $\lambda$.
Theorem 4.2.
Suppose the $l^{2}$-norm of the $k$-hop connectivity adjacency operator $\|\hat{A}_{k}(t)\|_{2}$ can be derived from Eq. (7). Then, as the $k$-hop approaches infinity ($k\to\infty$ ), the $\|\hat{A}_{k}(t)\|_{2}$ is bounded by the following inequality:
@@
The detailed proof of Eq. 8 is provided in Appendix B. To ensure numerical stability when computing powers of $\hat{A}_{k}(t)$, we establish the bounded condition: $\|\Phi_{k}(t)\|_{2}\leq 1$, which holds for all values of $k$ and $t$. In proposition 4.3, we demonstrate that $\Phi_{k}(t)$ exhibits smooth temporal evolution of the graph structure while maintaining a well-defined upper bound.
Proposition 4.3 (Convergence and Uniqueness).
Let $\Phi_{k}(t)$ be the normalized adjacency operator that satisfies the Lipschitz condition (Coddington & Levinson, 1955) $\|\Phi_{k}(t_{1})-\Phi_{k}(t_{2})\|_{2}\leq L\|t_{1}-t_{2}\|_{2}$ for some constant $L\geq 0$. Then, for a Lipschitz continuous activation function $\sigma$ and bounded weights $W_{k}^{(l)}$, the $k$-hop ODE-GCN admits a unique solution.
The proof from Picard’s existence theorem can be found in Appendix C.
Theorem 4.4 (Discretization of Age-Aware Continuous-Time Graph Convolution (Tang et al., 2024)).
Consider a series of brain networks $\{\mathcal{G}(t;\theta)\}_{t=0}^{T}$ parameterized by an age variable $\theta$. Let $Z(t)$ denote the embedding (or feature representation) of the STE-ODE network at time $t$. The continuous temporal evolution of the brain network can be expressed through the following age-modulated ODE-based graph convolution representation:
@@
where $H^{(l)}(t,\theta)$ representing the age-modulated layer-wise transformation at time $t$. Using a first-order approximation, the temporal dynamics described in Eq. 6 and theorem 4.4 can be discretized into the following form:
@@
$=Z(t)+\Delta t\sum_{l=1}^{L}\sigma\Big{(}\sum_{k=0}^{K-1}\Phi_{k}(t)X_{\theta}^{(l)}(t)W_{k}^{(l)}\Big{)},$
Here, $X_{\theta}^{(l)}(t)=X^{(l)}(t)\odot(\beta\cdot\theta)$ denotes the age-modulated feature matrix with learnable parameter $\beta\in(0,1)$ for age modulation strength at layer $l$, and $W_{k}^{(l)}$ represents the learnable weights associated with the $k$-th hop filter weight at layer $l$.
For $t\geq 1$, the temporal age-modulated graph embedding update is expressed as: